-8=-(x+4) [ditribute negative into the ()]
-8=-x-4
-8+4=-x-4+4 [add 4 on both sides]
-4=x
x=-4

Answer by ayelet(1)

-8=-(x+4) [ditribute negative into the parenthasis]
-8=-x-4
-8+4=-x-4+4 [add 4 on both sides]
-4=x
x=-4